OpenStudy (anonymous):

The equation, y = 12841(1.01)x , represents the average cost of an education at a private college, after 1990. What is the average cost in 1990 (x = 0)?

OpenStudy (bahrom7893):

12841

OpenStudy (anonymous):

y = 12841(1.01)^x y = 12841(1.01)^0 y = 12841(1) = $12841